a ball of mass m is dropped from the height h on a platform fixed at the top of a vertical spring. the platform is depressed by a distance x. what is the spring constant

Grade:12

2 Answers

Vikas TU
7 years ago
Eenrgy Conservation principle,
We would take the refernce at dropping point and let h be the height of the ball  from the platform
Thus,
0 + mgh = 0.5*k*x^2 + 0
k = 2mgh/x^2 is the answer.
Shashank
4 years ago
Let us assume that the ball is at a height 'h' from the ground.wkt U initial+KE initial=U final+KE final.initially it is at a height h.therefore mgh+0=1/2mv^2+0.we get v=√2gh.wkt s=ut+1/2at^2.u=0.therefore t=√2h/g.wkt F=ma.F=mv/t.put values for v and t and square both sides.we get F^2=(mg)^2.therefore answer=F=mg.
 
 
Hope it helps!!!!